OpenStudy (hba):

When n is divided by 12,the remainder is 6.What is the remainder when n is divided by 6 ?

OpenStudy (anonymous):

According to first statement : n = 12q + 6 According to second statement: n = 6p + r We are to find r..

OpenStudy (hba):

I am not very sure on how you generated the statements ?Please explain.

OpenStudy (anonymous):

See, there is formula : \[Dividend = Divisor \times Quotient + Remainder\]

OpenStudy (anonymous):

If I say 4 divides 12 leaving quotient as 3 and 0 : Then we can represent it as : \[12 = 4 \times 3 + 0\]

OpenStudy (anonymous):

and *remainder as 0..

OpenStudy (hba):

Ok then how is the Q(x) p and q

OpenStudy (anonymous):

I have assumed them as p and q in that two cases because we don't know them..

OpenStudy (anonymous):

@waterineyes what you given is true but there is two eqns and 3 unknowns... am I wrong?

OpenStudy (anonymous):

Yes we have 3 unknowns and this time I am remembering @mukushla because he is expert in finding 3 unknowns out of 1 equation given only.. Ha ha ha ha...

OpenStudy (tamtoan):

n = 12q + 6 = 6(2q +1) so n is a multiple of 6 ...:) when you that a multiple of 6 divide 6...is there a remainder ? :)

OpenStudy (anonymous):

when 6,,,...remainder should be 0 !!!!!

OpenStudy (anonymous):

tamtoan got the correct way..

OpenStudy (hba):

i have options so we can put them and check ?

OpenStudy (anonymous):

Sure..

OpenStudy (anonymous):

yep....

OpenStudy (anonymous):

18, 30, 42, 54 all leaves remainder as 6 when divided by 12.. So, you can easily see that they all are fully divided by 6, leaving remainder as 0..

OpenStudy (anonymous):

*leave

OpenStudy (anonymous):

Thanks to all........ I never think of bunchy equations but only facts... My bad.... i thought how it is possibl to solve it... my bad,... but thanks I got it

OpenStudy (hba):

0,1,2,3

OpenStudy (hba):

^ options ?

OpenStudy (hba):

Now you guys have pretty confused me lol

OpenStudy (hba):

@waterineyes Now tell me the math behind this :(

OpenStudy (anonymous):

I think @hba will be able to comprehend Ida behind it when you again check the first response of waterineyes

OpenStudy (anonymous):

Are you still not able to find the correct choice??

OpenStudy (anonymous):

followed by tamtoan's first response

OpenStudy (hba):

Well the answer is 0 but i could not understand the math behind it :(

OpenStudy (anonymous):

Till here you got : \[n = 12q + 6\]

OpenStudy (hba):

yes

OpenStudy (anonymous):

Now can you factor out 6 from that ??

OpenStudy (hba):

no.

OpenStudy (hba):

how to do that ?

OpenStudy (anonymous):

By using distributive property??

OpenStudy (anonymous):

\[ab + ac = a(b+c)\]

OpenStudy (hba):

n=6(2q+1)

OpenStudy (anonymous):

Yep. Now if I say that \(n = 3\times 4\) then it means that 3 and 4 are the factors of n.. Okay??

OpenStudy (hba):

ok

OpenStudy (hba):

so here the factors are 6 and 2q+1

OpenStudy (anonymous):

Good going..

OpenStudy (anonymous):

And you must remember that factors divide the number fully leaving NO REMAINDER BEHIND.. If \(n = 3 \times 4\) then 3 and 4 both will divide n fully leaving remainder as 0.. Getting or more explanation you want??

OpenStudy (hba):

more..

OpenStudy (anonymous):

See, here \(3 \times 4 = 12\) So here n = 12.. Now divide 12 by 3 first : Quotient - 4 Remainder - 0 Now divide 12 by 4: Quotient - 3 Remainder - 0 What did you get as remainder??

OpenStudy (anonymous):

It means that factors of any number will divide it fully leaving No Remainder.. It is obvious hba..

OpenStudy (anonymous):

More ??

OpenStudy (anonymous):

Still not getting tell me @hba

OpenStudy (hba):

gotcha:D

OpenStudy (hba):

So, 6/n=0 and 6/2q+1 = 0

OpenStudy (hba):

what next ?

OpenStudy (anonymous):

You have to find remainder no??

OpenStudy (anonymous):

That you have found I think... Why do you want to go next in this??

OpenStudy (anonymous):

Sorry, remainder of n/6 will be 0.. And not 6/n..

OpenStudy (anonymous):

Net got crushed??

OpenStudy (hba):

Sorry some issue with my internet connection,So the answer is 0 @waterineyes

OpenStudy (anonymous):

Yes it is...